# Quantum mechanics provides the physical basis of teleological evolutions

### v2

@@ 2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with the classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it.
@@ 2.5 Origin of the retrocausality
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality. Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept. In principle, classical logic should also go along with the coexistence of causality and retrocausality.. Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it. By the way, note that this is not an idle speculation, as might even seem. Indeed, we are dealing with a logical form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
@@ 2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it. Note that we never needed to look inside that unitary evolution $\hat{\mathbf{U}}_{1,2}$: see Tables I through IV of the Appendix. Therefore, this time-symmetrization does not only hold for quantum algorithms, it lends itself to an important generalization:
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule. The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5]. Note that this is what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
@@
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [15] might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
@@ 2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolution of the universe toward the value of life. We repeat here for convenience the reason for its possible quantum speedup and consequent teleological character.
Under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental constants, and each with its own value of the latter. Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”. The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life. Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life, in the other universes where this value is different there is absence of life.
We have seen in the Generalization subsection that any such building up of quantum correlation can always occur with quantum speedup in the sense that there is certainly the room for it. It can consequently be teleological in character.
@@
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life holds for the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from the former evolution by tracing all the rest. Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
@@
In summary, under the quantum cosmological assumption, the evolution of life on Earth can occur with quantum speedup and consequently be teleological in character; it might actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
@@
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would mean for the notion of the evolutions of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly. Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds. So why anticipate something designed for flight many millions of years in advance?
@@
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually, feathers turned out to be what was needed to conquer the air.
@@
All this is obviously destabilizing; it is a way of viewing the macroscopic reality that goes against the current (scientific) way of viewing it, and revamps the natural philosophy views that were ascribed by science to the unscientific beliefs. Of course, the fact that what occurs to us today is affected by the drying up of an area of Africa three millions of years ago is obvious, but that it is also affected by what we will be in three million years’ time, to the current way of thinking is more difficult to digest.
This also raises the question of how to describe a macroscopic evolution endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
Incidentally, it might also be interesting to derive the description of the evolution of life on Earth from that of the universe toward life by tracing out all the rest at the level of toy models of the two evolutions. Within the present theory, this might highlight significant characters of a possible teleological evolution of life on Earth.
By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter. However, naturally, classical physics turns out to work extremely well in describing macroscopic reality. Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions. This situation might point to a particular object of research. It could be interesting to study how the operation of “tracing-out all the rest” (at the toy model level) reconciles the teleological character of an evolution in the longer term with its classical character in the shorter term. This could also be a further perspective under which to address the problem of unifying classical and quantum theory.
@@
Note how the notion of the logical reversibility of computation played an essential role in the development of the notion of quantum computation. The present time-symmetric quantum description of quantum algorithms could be a further step in this direction. Involving a time-symmetric coexistence of the two possible time-directions of causality, it could be said that it exploits the objectively ideal time-reversibility of unitary quantum evolutions in the possibly deepest way.
@@ 3 Precedents
It is a work by Costa de Beauregard [24] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar. Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years. Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory, which also implied a form of retrocausality. See [25] for an account of this episode. It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [26, 27]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined. On the contrary, the present time-symmetric form of retrocausality is sort of universal for quantum correlations. The possible relation of those works to the present work, as things are now, would not be evident. Of course, it could be the object of study.
@@ 4 Conclusions
That the classical logic description of the quantum speedup must resort to retrocausality should be obvious. Indeed, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers. To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality. We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
@@
We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or, according to tastes, even right for them).
The other side of the coin of the destabilizing character of the work is, of course, its revolutionary character, always due to the retrocausality implicit in it. We should distinguish between what should be certain and what could be just plausible. Beginning with what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedup, which, forty years after its discovery, was still missing from mainstream quantum mechanics. (ii) It theoretically confirms the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has today in science and the scope of this notion, this should be the most revolutionary result of all. What is just plausible: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
Let us eventually note that all this would change our way of viewing physical reality. Curiously, it would be a case where the progress of science, the advent of quantum mechanics, brought us back to a way of viewing physical reality that preceded science.

### v3

@@ 2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with their classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it.
@@ 2.5 Origin of the retrocausality
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality. Associating it with a causality that always goes in the same time-direction, as customary, must be consequent to the de facto no-retrocausality precept. Classical logic should also go along with the coexistence of causality and retrocausality.. Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fundamental notion of quantum state superposition, the fact that the classical logic description of its generation necessarily involves retrocausality. By the way, note that this is not an idle speculation as it might even seem. Indeed, we are dealing with a form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
@@ 2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it. Note that we never needed to look inside the unitary evolution, for example $\hat{\mathbf{U}}_{1,2}$ of Tables I through IV of the Appendix. Thus, the need for this time-symmetrization does not only hold for quantum algorithms; there is the following generalization:
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
As shown in [5], the time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement. Note that this is exactly what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process equivalent in this case.
@@
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
By the way, if the only possible hidden variables are retrocausal in character as the ones we are dealing with, this dismissal can be ascribed to the already mentioned de facto rejection of retrocausality and finalism (the form it has in philosophy, different from the present trivial form).
@@ 2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life. This section expands on some facts that, for compactness, were just mentioned in the Introduction.
@@
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value becomes the possible teleological character of the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from the evolution of the universe toward life by tracing all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
@@
In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would be the room for this) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; it will be very destabilizing with respect to the present way of viewing the evolutions of the living.
@@
Let us begin with the evolution of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds. So why anticipate something designed for flight many millions of years in advance?
@@
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually feathers turned out to be what was needed to conquer the air.
@@
All this is obviously destabilizing; it is a way of viewing the time-development of macroscopic reality that goes against the current scientific (causal) way of viewing it. Indeed, it revamps the natural philosophy views that were ascribed by science to the category of the unscientific beliefs.
To the above example of teleological evolution of life, which concerned the evolution of the DNA, we should add another example that concerns the environment of the evolution of life. For example, what we humans are today has been affected by the drying up of an area of Africa in which our ancestors lived on trees three million years ago. Of course, there is nothing destabilizing in this. However, the time-symmetric fact that what we humans are today is also affected by what will concern our descendants in three million years’ time might be very difficult to digest. Indeed, it would profoundly change our (causal) way of viewing how things evolve in time.
Of course, all this is under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle and is thus hypothetical. What we would consider certain is the teleological character of quantum algorithms.
Things could be further explored by investigating, either theoretically or by means of toy models, the character of the “tracing over all the rest”. We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing over all the rest. We should look for a way of obtaining an evolution of the subsystem that is teleological in the longer term and (about) classical in the shorter term. If this were possible, of course, it would go in favor of the above quantum cosmological reading of that version of the Anthropic Principle. By the way, it might also be a way of addressing the problem of unifying quantum and classical theory.
Another question is how to describe a macroscopic evolution endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
Note the importance of the role that the notion of reversibility played in the development of quantum computation. The present work could be put on the same path. It could be said that the present time-symmetric form of causality exploits the objectively ideal time-reversibility of unitary quantum evolutions in the deepest possible way.
@@ 3 Precedents
It is a work by Costa de Beauregard [24, 25] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [26, 27]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, none of them concerns the quantum computational speedup. Their possible relation to the present work, as things are now, would not be evident.
@@ 4 Conclusions
That the classical logic description of the quantum speedup must resort to retrocausality should be obvious. Of course, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers. To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality. We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
@@
We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences.
The flip side of the destabilizing character of the work is its innovative character. Let us distinguish between what should be certain and what should be plausible. About what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics. (ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result. About what is just plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.